OpenStudy (anonymous):

5u + 3v = -16 2u + v = -8 Find the solution.

OpenStudy (anonymous):

\[5u = -16 - 3v\] \[u = \frac{ -16-3v }{ 5 }\] \[\frac{ -32-6v }{ 5 }+v = -8\] \[-32-6v=-40-5v\] \[-32+40=6v-5v\] \[v=8\] Then we replace v=8 in the first equation: \[5u=-16-24\] \[u=-40/5 \] \[u=-8\] So finally, \[ u=-8, v=8\]